# Learning when to rank: Estimation of partial rankings from sparse, noisy comparisons

### v1

@@ III.2 Real Match Datasets
Finally, we also compare the results inferred by our algorithm with those obtained by the BT model. Fig. 3(d) shows the number of nodes $N$ in each network as a function of the rankings $R$ inferred by the partial rankings algorithm, where the points are coloured according to the $\tau_{B}$ score between the rankings inferred by the PR and BT models. Again, we observe a general tendency for the $\tau_{B}$ scores to increase as $R$ approaches $N$.
@@ III.3 Case Study: CS Faculty Hiring
Figure 4: Partial rankings of CS departments according to faculty hiring patterns. (a) Barplot of BT strengths for the 205 PhD granting institutions included in the dataset, ordered according to their inferred BT rankings. Colours indicate the PR membership of each institution. A legend displaying the numerical values of these PR strengths is shown in panel (b). The names of the five institutions making up the strongest PR cluster and of the first five institutions in the second-strongest PR cluster (ordered in terms of their BT rank) are shown in the figure. (b) Barplot depicting the inverse of the USN ranking score for the 153 institutions for which data was available. Colours indicate the corresponding PR membership of each institution. Some notable inconsistencies between the USN rankings of some institutions and those inferred by pairwise comparison methods are shown in the figure.
@@
Figure 5: Partial rankings of CS departments and their external metadata. (a) BT rankings as a function of USN rankings for all 153 PhD granting institutions that had associated USN ranking. Points are coloured according to their PR rank. The black dashed line corresponds to the case $R_{BT}=R_{USN}$. (b) Average BT strength of a node’s out-neighbours as a function of the node’s out-degree. Points are coloured according to their PR rank. The horizontal lines represent the average BT score of the out-neighbours disaggregated by PR group. (c) Average endowment (in millions of USD) for the institutions in each partial rank. Error bars represent the standard deviations within each partial rank group. (d) Average endowment (in millions of USD) as a function of the USN rank, disaggregated by partial rankings. Error bars represent the standard deviations within each group of universities binned by USN rank (including ties). (e) Network representation of faculty hiring flows among the partial rank groups inferred by our algorithm. Each node represents a partial ranking, with labels and colours indicating the corresponding group. Node sizes are proportional to the number of institutions assigned to each partial ranking. Directed edges represent the hiring flows between groups, with edge thickness proportional to the net flow. Edge colours transition from the origin node’s colour at the base to the destination node’s colour at the arrowhead and the net flow value is annotated on each edge.
@@
As a case study, we apply our algorithm to a dataset of faculty hires in Computer Science (CS) departments at U.S. higher education institutions, compiled in clauset2015systematic for analysis of hierarchies in faculty hiring using a different ranking methodology. Each node represents one of the 205 PhD-granting institutions included in the study and a directed edge $(i,j)$ exists from institution $i$ to institution $j$ for each individual who earned their doctorate at $i$ and held a tenure or tenure-track faculty position at $j$ during the collection period from May 2011 to March 2012. For 153 of these institutions, their U.S. News & World Report rankings (USN rankings) are available in the form of node metadata, which we use to compare against the rankings and partial rankings inferred by the two pairwise ranking methods employed in this study. Additionally, we manually collected endowment data for 151 of the 153 institutions with associated USN rankings.
@@
Fig. 4(a) shows the results of applying the Bradley-Terry model and our partial rankings algorithm to the entire set of 205 PhD granting institutions. Each bar represents an institution, ordered by its BT rank, with bar heights representing the BT strength scores. Bars are colour-coded according to their PR rankings, with corresponding strengths shown in the legend in Fig. 4(b). The rankings are dominated by a small group of elite universities displaying significantly higher strengths, followed by a sharp decline in strength values further down the list. The PR and BT rankings show strong ordinal association ($\tau_{B}=0.82$; p-value $=1.2\times 10^{-53}$) with only a small number of rank violations —instances where $\text{rank}_{BT}(A)>\text{rank}_{BT}(B)$ but $\text{rank}_{PR}(A)<\text{rank}_{PR}(B)$— at the boundary between the two lowest-scoring PR groups, possibly indicating a lack of information to accurately determine the group boundaries.
@@
In Fig. 4(b), we compare the partial rankings inferred by our model with the U.S. News & World Report rankings. Each bar in the plot represents a PhD-granting institution, ordered by its USN rank, with bar heights corresponding to the inverse of their USN rank. Again, bars are colour-coded according to their partial rankings inferred by the PR model.
@@
One first thing to notice with regards to the USN rankings is that they allow for partial rankings. For example, the four highest-scoring universities according to the USN methodology are all assigned an equal rank of 1. Fig. 4(b) reveals that institutions with equal USN ranking often belong to the same partial rankings according to the PR model. However, the PR model generally produces broader rankings, grouping institutions that span a wider range of USN ranks. Another striking feature is the significantly higher number of rank violations, although the overall ordinal association remains strong ($\tau_{B}=0.76$; p-value $=1.37\times 10^{-33}$). Some of these rank violations are particularly striking. For example, Harvard University, ranked $5^{th}$ in the BT model and belonging to the top-scoring PR group, is ranked $17^{th}$ according to the USN rankings, behind institutions including the University of Maryland, College Park ($26^{th}$ in the BT model and in the third PR group) that scored significantly lower according to the hiring network structure. Similarly, the University of Chicago is ranked $19^{th}$ by the BT model and belongs to the second highest-scoring PR group but is ranked $35^{th}$ in the USN rankings. While the U.S. News & World Report utilises a wide array of metrics to conduct its rankings, such as peer assessment scores, graduation and retention rates, and financial resources, this result would appear to suggest that these factors may not directly align with academic job market preferences. In the case of Harvard University or the University of Chicago, one might surmise that the prestigious reputation of the institutions might be influencing the job market more than their USN rankings. The opposite effect can also observed. For instance, Northeastern University, ranked $61^{st}$ in the USN rankings, is assigned $178^{th}$ place by the BT model and belongs to the lowest PR group despite its considerably higher USN ranking. This suggests that perhaps this program was undervalued in the job market relative to its academic credentials during the period studied.
@@
To evaluate how the BT rankings compare to the USN rankings, Fig. 5(a) plots the BT rank of each institution against its corresponding USN rank, with points colour-coded based on their PR rank. The highest-ranking institutions, according to both models, tend to concentrate around the $R_{BT}=R_{USN}$ line, where $R_{BT}$ and $R_{USN}$ represent the Bradley-Terry and USN rankings respectively —indicating close agreement between the two ranking systems for top-tier institutions. As we move towards lower-ranked institutions, we observe an increasing divergence between the rankings. While the overall ordinal association remains strong ($\tau_{B}=0.73$, p-value $=1.07\times 10^{-38}$), the spread in rankings widens, and institutions with a given USN rank are often assigned considerably different BT ranks, with the BT model frequently assigning these institutions higher positions compared to the USN rankings. This discrepancy is particularly pronounced among institutions in the two lowest-scoring PR groups.
@@
An interesting question about the pairwise ranking methods considered in this work is whether an institution’s rank benefits more from “who you win against” than from the sheer number of “wins”. In other words, does an institution’s ranking improve more by producing a large number of doctoral graduates who secure tenure-track positions across a wide range of institutions or by producing fewer graduates who secure positions at highly ranked institutions? To address this, in Fig. 5(b), we plot the average BT strength of an institution’s out-neighbors (capturing the quality of an institution’s “wins”) as a function of its out-degree (representing the total number of “wins” an institution has). We observe a clear positive correlation (Pearson $p=0.78$; p-value $=1.32\times 10^{-43}$) with the highest ranked institutions having both a large number of “wins” (i.e. producing a large number of graduates that go on to secure tenure-track positions at other institutions), as well as out-neighbours with high BT strengths (i.e. their graduates tend to secure positions at high-ranking institutions). Perhaps more revealing is the fact that the PR groups appear to stratify according to the average BT strength of their out-neighbours, suggesting a well-defined hierarchy characterised by a predominance in horizontal mobility as opposed to vertical mobility. This observation is reinforced by analysing the network of mobility flows between the PR groups; see Fig. 5(e). The results reveal a well-defined hierarchy with minimal upward mobility across ranks. Most mobility occurs downward within the hierarchy, a trend partly explained by the relatively small size of the top-ranked groups. Nonetheless, horizontal mobility emerges as a critical driver of hiring dynamics across all levels, particularly within the lower-ranked groups, where it becomes the dominant hiring pattern. Horizontal mobility is particularly pronounced at the topmost level of the hierarchy, where the top five institutions comprising this elite group hire over twice as many faculty members from within their group as they do from all other PhD granting institutions combined.
@@ IV Conclusion
We have introduced a probabilistic generative model for inferring partial rankings in directed networks and a fast, fully nonparametric, agglomerative algorithm for efficient inference. We have shown that, particularly with limited observations available, our model can provide a more parsimonious description of pairwise comparison data than models that inherently assume complete rankings such as the BT model. Specifically, in extremely sparse regimes, our model effectively recognises that insufficient evidence exists to infer any meaningful ranking among the compared entities. As the network’s connectivity increases, partial rankings typically emerge as a more compact and accurate description of the data. Finally, as we move towards more dense regimes, we reach a point where sufficient information is available to infer a complete ordinal ranking of all the nodes. When applied to a network of faculty hiring among U.S. computer science departments, our algorithm inferred five distinct partial rankings, which align closely with the ordinal rankings inferred by the BT model and reveal a well-defined hierarchy dominated by a small number of elite universities. The inferred rankings also highlight the limited upward mobility within the hierarchy, with lateral and downward movements being more prevalent.
@@
There are several directions in which our work can be extended. While widely used, the classical Bradley-Terry model represents a relatively simplistic approach to pairwise interactions between entities. Numerous extensions to the model have been proposed, including accounting for ties rao1967ties ; davidson1970extending , home-field advantage agresti2012categorical , randomness in match outcomes, and imbalances in strength or skill between the average pair of players jerdee2024luck . Any of these extensions could, in principle, be incorporated into our partial rankings framework, enabling the development of more expressive models.
@@
Another characteristic of both the Bradley-Terry model and our partial rankings framework is that they do not explicitly account for the placement of edges in the network—that is, the likelihood does not include a term that models the probability of observing a given set of matches in the first place. As we observed in Sec. III.3, similarly (partially) ranked institutions tend to interact primarily with one another. A similar observation holds for the ATP tennis dataset, where the strongest players, who advance through the tournament, tend to play more matches and preferentially against each other. While other methods explicitly model edge placement de2018physical ; peixoto2022ordered , this aspect remains absent from the current Bradley-Terry and partial rankings models. Incorporating it into the framework could help to explore how network topology influences ranking recovery.
@@
Finally, our framework could be extended to consider other cases of interest, such as dynamic rankings, where the ranks of the individual entities can rise or fall over time della2024model , personalised rankings, in which ranking results are tailored to individual users based on their preferences rendle2012bpr , or a combination of both.
@@ Acknowledgments
The authors acknowledge funding support from the HKU-100 Start Up Fund and an HKU Urban Systems Institute Fellowship Grant, and Max Jerdee for helpful discussions about the datasets.
Appendix A Time Complexity

### v2

@@ III.2 Real Match Datasets
Finally, we also compare the results inferred by our algorithm with those obtained by the BT model. Fig. 3(d) shows the number of nodes $N$ in each network as a function of the rankings $R$ inferred by the partial rankings algorithm, where the points are colored according to the $\tau_{B}$ score between the rankings inferred by the PR and BT models. Again, we observe a general tendency for the $\tau_{B}$ scores to increase as $R$ approaches $N$.
@@ III.3 Case Study: CS Faculty Hiring
Figure 4: Partial rankings of CS departments according to faculty hiring patterns. (a) Barplot of BT strengths for the 205 PhD-granting institutions included in the dataset, ordered according to their inferred BT rankings. Colors indicate the PR membership of each institution. A legend displaying the numerical values of these PR strengths is shown in panel (b). The names of the five institutions making up the strongest PR cluster and of the first five institutions in the second-strongest PR cluster (ordered in terms of their BT rank) are shown in the figure. (b) Barplot depicting the inverse of the USN ranking score for the 153 institutions for which data was available. Colors indicate the corresponding PR membership of each institution. Some notable inconsistencies between the USN rankings of some institutions and those inferred by pairwise comparison methods are shown in the figure.
@@
Figure 5: Partial rankings of CS departments and their external metadata. (a) BT rankings as a function of USN rankings for all 153 PhD-granting institutions that had associated USN ranking. Points are colored according to their PR rank. The black dashed line corresponds to the case $R_{BT}=R_{USN}$. (b) Average BT strength of a node’s out-neighbors as a function of the node’s out-degree. Points are colored according to their PR rank. The horizontal lines represent the average BT score of the out-neighbors disaggregated by PR group. (c) Average endowment (in millions of USD) for the institutions in each partial rank. Error bars represent the standard deviations within each partial rank group. (d) Average endowment (in millions of USD) as a function of the USN rank, disaggregated by partial rankings. Error bars represent the standard deviations within each group of universities binned by USN rank (including ties). (e) Network representation of faculty hiring flows among the partial rank groups inferred by our algorithm. Each node represents a partial ranking, with labels and colors indicating the corresponding group. Node sizes are proportional to the number of institutions assigned to each partial ranking. Directed edges represent the hiring flows between groups, with edge thickness proportional to the net flow. Edge colors transition from the origin node’s color at the base to the destination node’s color at the arrowhead and the net flow value is annotated on each edge.
@@
As a case study, we apply our algorithm to a dataset of faculty hires in Computer Science (CS) departments at U.S. higher education institutions, originally compiled in clauset2015systematic for analyzing hierarchies in academic hiring using a different ranking methodology. Each node in the network represents one of the 205 PhD-granting institutions included in the study, and a directed edge $(i,j)$ is included for each individual who received their doctorate from institution $i$ and held a tenure or tenure-track faculty position at institution $j$ during the data collection period (May 2011 to March 2012). For 153 of these institutions, we have access to their U.S. News & World Report (USN) rankings, which we use as node metadata to compare against the rankings and partial rankings inferred by the two pairwise ranking methods explored in this study. The USN rankings are based on a broad set of criteria—including reputation, research output, student selectivity, and resources—and, as such, are not expected to directly reflect hiring dynamics. However, it is nonetheless useful to examine the extent to which these widely cited rankings align with those inferred from the academic job market dynamics. Additionally, we manually collected endowment data for 151 of the 153 institutions with USN rankings to explore potential correlations with inferred prestige.
@@
Fig. 4(a) shows the results of applying the BT model and our PR algorithm to the entire set of 205 PhD-granting institutions. Each bar represents an institution, ordered by its BT rank, with bar heights representing the BT strength scores. Bars are color-coded according to their PR rankings, with corresponding strengths shown in the legend in Fig. 4(b). The rankings are dominated by a small group of elite universities displaying significantly higher strengths, followed by a sharp decline in strength values further down the list. The PR and BT rankings show strong ordinal association ($\tau_{B}=0.82$; p-value $=1.2\times 10^{-53}$) with only a small number of rank violations—instances where $\text{rank}_{BT}(A)>\text{rank}_{BT}(B)$ but $\text{rank}_{PR}(A)<\text{rank}_{PR}(B)$— at the boundary between the two lowest-scoring PR groups, possibly indicating a lack of information to accurately determine the group boundaries.
@@
In Fig. 4(b), we compare the partial rankings inferred by our model with the U.S. News & World Report rankings. Each bar in the plot represents a PhD-granting institution, ordered by its USN rank, with bar heights corresponding to the inverse of their USN rank. Again, bars are color-coded according to their partial rankings inferred by the PR model.
@@
One first thing to notice with regards to the USN rankings is that they allow for partial rankings. For example, the four highest-scoring universities according to the USN methodology are all assigned an equal rank of 1. Fig. 4(b) reveals that institutions with equal USN ranking often belong to the same partial rankings according to the PR model. However, the PR model generally produces broader rankings, grouping institutions that span a wider range of USN ranks. Another striking feature is the significantly higher number of rank violations, although the overall ordinal association remains strong ($\tau_{B}=0.76$; p-value $=1.37\times 10^{-33}$). Some of these rank violations are particularly striking. For example, Harvard University, ranked $5^{th}$ in the BT model and belonging to the top-scoring PR group, is ranked $17^{th}$ according to the USN rankings, behind institutions including the University of Maryland, College Park ($26^{th}$ in the BT model and in the third PR group) that scored significantly lower according to the hiring network structure. Similarly, the University of Chicago is ranked $19^{th}$ by the BT model and belongs to the second highest-scoring PR group but is ranked $35^{th}$ in the USN rankings. While the U.S. News & World Report utilizes a wide array of metrics to conduct its rankings, such as peer assessment scores, graduation and retention rates, and financial resources, this result would appear to suggest that these factors may not directly align with academic job market preferences. In the case of Harvard University or the University of Chicago, one might surmise that the prestigious reputation of the institutions might be influencing the job market more than their USN rankings. The opposite effect can also be observed. For instance, Northeastern University, ranked $61^{st}$ in the USN rankings, is assigned $178^{th}$ place by the BT model and belongs to the lowest PR group despite its considerably higher USN ranking. This suggests that perhaps this program was undervalued in the job market relative to its academic credentials during the period studied.
@@
To evaluate how the BT rankings compare to the USN rankings, Fig. 5(a) plots the BT rank of each institution against its corresponding USN rank, with points color-coded based on their PR rank. The highest-ranking institutions, according to both models, tend to concentrate around the $R_{BT}=R_{USN}$ line, where $R_{BT}$ and $R_{USN}$ represent the BT and USN rankings respectively—indicating close agreement between the two ranking systems for top-tier institutions. As we move towards lower-ranked institutions, we observe an increasing divergence between the rankings. While the overall ordinal association remains strong ($\tau_{B}=0.73$, p-value $=1.07\times 10^{-38}$), the spread in rankings widens, and institutions with a given USN rank are often assigned considerably different BT ranks, with the BT model frequently assigning these institutions higher positions compared to the USN rankings. This discrepancy is particularly pronounced among institutions in the two lowest-scoring PR groups.
@@
An interesting question about the pairwise ranking methods considered in this work is whether an institution’s rank benefits more from “who you win against” than from the sheer number of “wins”. In other words, does an institution’s ranking improve more by producing a large number of doctoral graduates who secure tenure-track positions across a wide range of institutions or by producing fewer graduates who secure positions at highly ranked institutions? To address this, in Fig. 5(b), we plot the average BT strength of an institution’s out-neighbors (capturing the quality of an institution’s “wins”) as a function of its out-degree (representing the total number of “wins” an institution has). We observe a clear positive correlation (Pearson $p=0.78$; p-value $=1.32\times 10^{-43}$) with the highest ranked institutions having both a large number of “wins” (i.e. producing a large number of graduates that go on to secure tenure-track positions at other institutions), as well as out-neighbors with high BT strengths (i.e. their graduates tend to secure positions at high-ranking institutions). Perhaps more revealing is the fact that the PR groups appear to stratify according to the average BT strength of their out-neighbors, suggesting a well-defined hierarchy characterized by a predominance in horizontal mobility as opposed to vertical mobility. This observation is reinforced by analyzing the network of mobility flows between the PR groups; see Fig. 5(e). The results reveal a well-defined hierarchy with minimal upward mobility across ranks. Most mobility occurs downward within the hierarchy, a trend partly explained by the relatively small size of the top-ranked groups. Nonetheless, horizontal mobility emerges as a critical driver of hiring dynamics across all levels, particularly within the lower-ranked groups, where it becomes the dominant hiring pattern. Horizontal mobility is also particularly pronounced at the topmost level of the hierarchy, where the top five institutions comprising this elite group hire over twice as many faculty members from within their group as they do from all other PhD-granting institutions combined.
@@ IV Conclusion
We have introduced a probabilistic generative model for inferring partial rankings in directed networks and a fast, fully nonparametric, agglomerative algorithm for efficient inference. We have shown that, particularly with limited observations available, our model can provide a more parsimonious description of pairwise comparison data than models that inherently assume complete rankings, such as the BT model. Specifically, in extremely sparse regimes, our model effectively recognizes that insufficient evidence exists to infer any meaningful ranking among the compared entities. As the network’s connectivity increases, partial rankings typically emerge as a more compact and accurate description of the data. Finally, as we move towards more dense regimes, we reach a point where sufficient information is available to infer a complete ordinal ranking of all the nodes. When applied to a network of faculty hiring among U.S. computer science departments, our algorithm inferred five distinct partial rankings, which align closely with the ordinal rankings inferred by the BT model and reveal a well-defined hierarchy dominated by a small number of elite universities. The inferred rankings also highlight the limited upward mobility within the hierarchy, with lateral and downward movements being more prevalent.
@@
There are several directions in which our work can be extended. While widely used, the classical BT model represents a relatively simplistic approach to pairwise interactions between entities. Numerous extensions to the model have been proposed, including accounting for ties rao1967ties ; davidson1970extending , home-field advantage agresti2012categorical , randomness in match outcomes, and imbalances in strength or skill between the average pair of players jerdee2024luck . Any of these extensions could, in principle, be incorporated into our partial rankings framework, enabling the development of more expressive models.
@@
Another characteristic of both the BT model and our partial rankings framework is that they do not explicitly account for the placement of edges in the network—that is, the likelihood does not include a term that models the probability of observing a given set of matches in the first place. As we observed in Sec. III.3, similarly (partially) ranked institutions tend to interact primarily with one another. A similar observation holds for the ATP tennis dataset, where the strongest players, who advance through the tournament, tend to play more matches and preferentially against each other. While other methods explicitly model edge placement de2018physical ; peixoto2022ordered , this aspect remains absent from the current BT and partial rankings models. Incorporating it into the framework could help to explore how network topology influences ranking recovery.
@@
Our framework could also be extended to consider other cases of interest, such as dynamic rankings, where the ranks of the individual entities can rise or fall over time della2024model , personalised rankings, in which ranking results are tailored to individual users based on their preferences rendle2012bpr , or a combination of both.
Finally, while we have developed a fully Bayesian nonparametric framework to infer partial rankings, we have focused on inferring point estimates of these rankings without considering uncertainty quantification as done in other works pearce2024bayesian ; peixoto2022ordered . However, as noted in Sec. II.3, the Bayesian nature of our model lends itself naturally to the exploration of the full posterior, which can be explored via the development of suitable inference algorithms—such as Expectation-Maximization (EM) or Markov Chain Monte Carlo (MCMC) methods—to enable posterior sampling or approximation, thereby allowing for principled uncertainty quantification in inferred rankings.
@@ Acknowledgments
The authors thank Max Jerdee for useful discussions about the datasets, and Michael Pearce and Elena Erosheva for helpful suggestions to improve the initial manuscript. The authors acknowledge funding support from the HKU-100 Start Up Fund and an HKU Urban Systems Institute Fellowship Grant.
Appendix A Minimization of $f(R)$
We prove that the function
$f(R)=\log\binom{N-1}{R-1}+\log\binom{N}{n_{1}\ldots n_{R}}$ (39)
has a global minimum at $R=1$.
We begin proving that the second term in Eq. 39 is maximized when all groups are of equal size such that $n_{r}=n=N/R~{}~{}\forall r$.
Proposition 1.
Let $N\in\mathbb{Z}^{+}$ and let $\{n_{r}\}\in\mathbb{Z}^{+}$ be a set of positive integers such that $\sum_{r=1}^{R}n_{r}=N$. Then the multinomial coefficient $\binom{N}{n_{1}\ldots n_{R}}$ is maximized when $n_{r}=n=N/R~{}~{}\forall r$.
Proof.
By definition
$\binom{N}{n_{1}\ldots n_{R}}=\frac{N!}{n_{1}!\ldots n_{R}!}.$ (40)
Using Stirling’s approximation of the factorial,
$\log k!\approx k\log k-k+\frac{1}{2}\log(2\pi k),$ (41)
we have that, to first order
$\binom{N}{n_{1}\ldots n_{R}}$ $=N\log N-N-\sum_{r=1}^{R}(n_{r}\log n_{r}-n_{r})$ (42)
$=N\log N-\sum_{r=1}^{R}n_{r}\log n_{r},$ (43)
where we have used the fact that $\sum_{r}n_{r}=N$.
To maximize Eq. 40, we then have to solve the following constrained optimization problem
$\max_{\{n_{r}\}}\left(N\log N-\sum_{r=1}^{R}n_{r}\log n_{r}\right)$ (44)
[…]
which corresponds to extremizing the following Lagrangian
($\mathcal{L}(\{n_{r}\},\lambda)=-\sum_{r=1}^{R}n_{r}\log n_{r}+\lambda\left(\sum_{r=1}^{R}n_{r}-N\right),$ (46)
where we have ignored constant terms. Taking the partial derivatives with respect to $n_{r}$ and equating to zero, we have
$\frac{\partial\mathcal{L}}{\partial n_{r}}=-(\log n_{r}+1)+\lambda=0\implies n_{r}=e^{\lambda-1}\equiv C.$ (47)
From the sum constraint, we then have
$\sum_{r=1}^{R}n_{r}=RC=N\implies C=n_{r}=\frac{N}{R}~{}~{}\forall r.$ (48)
Given Proposition 1 and in view of the monotonicity of the logarithm, it follows that the second term on the right-hand side of Eq. 39 is maximized when all groups are of equal size.
Given this result, we can now prove our original statement.
Proposition 2.
The function
$f(R)=\log\binom{N-1}{R-1}+\log\binom{N}{n_{1}\ldots n_{R}}$
has a global minimum at $R=1$ with $f(R=1)=0$.
Proof.
From Proposition 1 and by virtue of the monotonicity of the logarithm, it follows that Eq. 39 is upper bounded by
$f(R)\leq\log\binom{N-1}{R-1}+\log\frac{N!}{n!^{R}},$ (49)
with $n=N/R$. Eq. 49 can be written as
$f(R)$ $\leq\log\binom{N-1}{R-1}+\log N!-R\log n!$ (50)
$\approx\log\binom{N-1}{R-1}+N\log N-N-R\left(\frac{N}{R}\log\frac{N}{R}-\frac{N}{R}\right)$ (51)
$=\log\binom{N-1}{R-1}+N\log R,$ (52)
where we have once more made use of Stirling’s approximation. The first term on the right-hand side of Eq. 52 is a concave function of $R$, reaching its minimum value of zero at both $R=1$ and $R=N$. The second term, however, increases monotonically with $R$, starting from a minimum of zero when $R=1$. Since these two terms combine additively in Eq. 52, their sum must attain its global minimum of zero at $R=1$, where both terms simultaneously achieve their lowest values. Finally, since the binomial and multinomial coefficients are always greater than or equal to one, it follows that Eq 39 is also lower bounded by zero, meaning it must attain a global minimum of zero at $R=1$. ∎
Appendix B Time Complexity